What is the energy needed to take one neutron or o
Answer #1

a) mass of Nytrogen = 14.0067 u, mass of neutron = 1.008665 u, mass of proton = 1.007276 u. mass of carbon = 13.003355

removal of proton

14N 7 ---- 13C6 + p

14.0067 ---------13.003355+ 1.007276 = 14.010631 u

energy required to remove proton = (14.0067 -  14.010631 )931.5= 3.66 MeV apprx

b) to remove neutron,

14N7 -----13N7 + N

mass of 13N7 ( isotope oof Nitrogen) = 13.0057386

14.0067 = 13.0057386 + 1.008665 = 14.0144036

energy required to remove neutron= (14.0144036 - 14.0067)931.5 = 7.1759 mev apprx

energy required to remove neutron is more